# Non-planar ends are continuously unforgettable

### v1

@@ 1 Introduction
A fundamental tool in studying mapping class groups of finite-type surfaces is the forgetful homomorphism: Given a non-compact finite-type surface $S$ and a point $p\in S$, there is a surjective homomorphism $\operatorname{PMCG}(S\smallsetminus\{p\})\to\operatorname{PMCG}(S)$ obtained by “forgetting the puncture $p$”, where $\operatorname{PMCG}(S)$ is the pure mapping class group. For the forgetful homomorphism, the surface in the codomain has one less end than the surface in the domain. The motivating question of this article is to ask if the same phenomenon can occur in the setting of infinite-genus surfaces without planar ends, or in other words,
@@
Let $M$ and $M^{\prime}$ be orientable 2-manifolds, and suppose $M$ has infinite genus and has no planar ends. If $G$ is a mostly pure large subgroup of $\operatorname{MCG}(M)$ and $G^{\prime}$ is a large subgroup of $\operatorname{MCG}(M^{\prime})$, then every continuous epimorphism $G\to G^{\prime}$ is induced by an homeomorphism.
@@
Next, recall that an object $A$ in a category is Hopfian if every epimorphism $A\to A$ is an automorphism. It is a standard fact that every finitely generated residually finite group is Hopfian (in the category of abstract groups). As the mapping class group of a finite-type surface is residually finite (see [14, §6.4]) and finitely generated (see [14, §4.3]), every finitely generated subgroup of the mapping class group of a finite-type orientable 2-manifold is Hopfian. One of the motivating problems for this article is to understand when $\operatorname{MCG}(M)$ and $\operatorname{PMCG}(M)$ are Hopfian groups. As a corollary of Theorem 5.1, we establish a class of big mapping class groups that are Hopfian in the category of topological groups, see Problem 4.5 of [1].
@@
A 2-manifold $M$ is perfectly self-similar if $M\#M$ is homeomorphic to $M$ and for every proper compact subset $K$ of $M$ there exists $f\in\operatorname{Homeo}(M)$ such that $f(K)\cap K=\varnothing$.
@@
And as a corollary, we have the Hopfian property.
@@ Acknowledgements
The authors thank Greg Conner for a helpful discussion about infinitely multiplicative homomorphisms.
@@ 2.2 Mapping class groups and their topology
It will be useful to have a more workable description of this topology; this involves using a corollary to the Alexander method (see [14, §2.3] for finite-type surfaces and see [15] or [19] for infinite-type surfaces).
@@ 2.3 Twists
We already established the backwards direction above, so we only need to show the forwards direction. We will argue the contrapositive: let $A$ fail to be locally finite. Then there exists a curve $c$ and a subsequence of $\{a_{n_{j}}\}$ such that $i(c,a_{n_{j}})\neq 0$ for all $j\in\mathbb{N}$. Let $T_{m}=\prod_{j=1}^{n_{m}}t_{a_{j}}^{k_{j}}$, and let $F_{m}=T_{m}^{-1}T_{m+1}$. Note that if $\lim_{n\to\infty}\prod_{j=1}^{n}t_{a_{n}}^{k_{n}}$ exists, then the sequence $\{T_{m}\}_{m\in\mathbb{N}}$ converges, and hence $\{F_{m}\}_{m\in\mathbb{N}}$ converges to the identity. We will show that the sequence $\{F_{m}\}$ fails to converge to the identity, which will finish the proof.
Observe that $F_{m}(c)=t_{a_{n_{m+1}}}^{k_{n_{m+1}}}(c)$. For each $m\in\mathbb{N}$, choose a curve $b_{m}$ satisfying $i(c,b_{m})=0$ and $i(a_{n_{m+1}},b_{m})>0$. Then, by [14, Proposition 3.4],
$i(F_{m}(c),b_{m})=i(t_{a_{n_{m+1}}}^{k_{n_{m+1}}}(c),b_{m})\geq|k_{n_{m+1}}|i(a_{n_{m+1}},c)i(a_{n_{m+1}},b_{m})-i(c,b_{m}),$
and hence $i(F_{m}(c),b_{m})>0$, as $i(c,b_{m})=0$ and all the other terms on the right are positive. This implies that $F_{m}(c)\neq c$ for all $m\in\mathbb{N}$, and in particular, there exists a neighborhood of the identity in $\operatorname{MCG}(S)$ that the sequence $\{F_{m}\}_{m\in\mathbb{N}}$ never enters, namely $\{f\in\operatorname{MCG}(S):f(c)=c\}$. Therefore, the sequence $\{F_{m}\}$ does not converge. ∎
@@ 3 Generating compactly supported mapping classes
Let us begin by recalling the Humphries generating set (see [14, §4.4.3 and §4.4.4]). Humphries’s theorem tells us that the Dehn twists about the curves shown in Figure 1 generate $\operatorname{MCG}(S,\partial S)$; in particular, for a compact surface with nonempty boundary, the Humphries generating set consists of $2g+b$ non-separating Dehn twists (the figure shows a genus three surface with three boundary components, but it should be clear how to generalize to higher genus and more boundary components). We note that Humphries’s theorem requires the surface to be orientable and nonplanar. We will build on Humphries’s theorem to construct a generating set for $\Gamma_{c}(M)$ when $M$ is an orientable infinite-genus 2-manifold without planar ends.
@@
In what follows, we will require a generating set for the mapping class group of a four-holed sphere. Let $R$ be a compact four-holed sphere. If $a$ and $b$ are curves with $i(a,b)=2$, then $t_{a}$ and $t_{b}$ together with the Dehn twists along any three of the four boundary components of $R$ generate $\operatorname{MCG}(R,\partial R)$ (this is an exercise combining the computation of generators for the mapping class group of a pair of pants [14, §3.6.4] and a four-punctured sphere [14, §4.2.4], the inclusion homomorphism [14, Proposition 3.18], and the lantern relation [14, Proposition 5.1]).
Given a compact subsurface $\Sigma$ of a 2-manifold $M$ such that each component of $\partial\Sigma$ is separating and essential, the inclusion of $\Sigma$ into $M$ induces a monomorphism $\operatorname{MCG}(\Sigma,\partial\Sigma)\to\operatorname{MCG}(M)$, see [14, Theorem 3.18]. Abusing notation, we will identify $\operatorname{MCG}(\Sigma,\partial\Sigma)$ with its image.
@@
Let $\{\Sigma_{n}\}_{n\in\mathbb{N}}$ be the compact exhaustion of $M$ given by Lemma 3.1. Then $\Gamma_{c}(M)=\bigcup_{n\in\mathbb{N}}\operatorname{MCG}(\Sigma_{n},\partial\Sigma_{n})$. We therefore need to give an Alexander chain $\mathcal{A}$ such that the Dehn twists about the curves in $\mathcal{A}$ contained in $\Sigma_{n}$ generate $\operatorname{MCG}(\Sigma_{n},\partial\Sigma_{n})$. Let us construct $\mathcal{A}$ and then argue that it satisfies this property.
@@
Let $S_{0}=\Sigma_{0}$, and for $n\in\mathbb{N}$, let $S_{n}$ be the closure of $\Sigma_{n}\smallsetminus\Sigma_{n-1}$. Let $\mathcal{A}_{0}$ consist of any two simple closed curves with geometric intersection number one; hence, the curves in $\mathcal{A}_{0}$ generate $\operatorname{MCG}(\Sigma_{0},\partial\Sigma_{0})$ (see [14, §3.6.4]). For $i\in\{2,3\}$, let $F_{i}$ be the genus one compact surface with $i$ boundary components. Let $C_{i}$ be the Alexander chain in $F_{i}$ shown in Figure 2. For $n\in\mathbb{N}$, let $i(n)\in\{2,3\}$ such that $S_{n}$ is homeomorphic to $F_{i(n)}$, and fix a homeomorphism $f_{n}\colon\thinspace F_{i(n)}\to S_{n}$. Let $\mathcal{A}^{\prime}$ be the chain defined by $\mathcal{A}^{\prime}=\mathcal{A}_{0}\cup\bigcup_{n=1}^{\infty}f_{n}(C_{i(n)})$.
We now add curves to the collection $\mathcal{A}^{\prime}$ so that it forms an Alexander chain. Each of the curves in Figure 2 are colored either blue or red; each surface in the figure contains a single red curve that intersects each of the remaining curves—the blue curves—exactly once. Note that, in each surface, the blue curves give a pants decomposition. Designate one curve in $\mathcal{A}_{0}$ to be blue. Now, let $\mathcal{A}^{\prime}_{b}$ denote the subset of $\mathcal{A}^{\prime}$ containing all the blue curves. Each component of $M\smallsetminus\mathcal{A}^{\prime}_{b}$ is a four-holed sphere. Enumerate the components as follows: define $\Omega_{n}$ to be the unique component that is contained in $\Sigma_{n}$ but not in $\Sigma_{n-1}$. Choose a non-separating curve $a_{n}$ in $M$ that is contained in $\Omega_{n}$ and such that $\mathcal{A}^{\prime}\cup\{a_{n}\}$ is a chain. Let $\mathcal{A}=\mathcal{A}^{\prime}\cup\{a_{n}:n\in\mathbb{N}\}$. By construction, $\mathcal{A}$ is an Alexander chain.
@@ 4 Irreducible epimorphisms
If $g\in\Gamma_{c}(M)$ is nontrivial, then there exists a finite-type subsurface $\Sigma\subset M$ containing the support of $g$ such that each boundary component of $\Sigma$ is an essential simple closed curve. In this case, the inclusion $\Sigma\hookrightarrow M$ induces a monomorphism $\operatorname{MCG}(\Sigma,\partial\Sigma)\to\Gamma_{c}(M)$ whose image contains $g$. Therefore, we can view $g$ as an element of $\operatorname{MCG}(\Sigma,\partial\Sigma)$, which is torsion free (see [14, Corollary 7.3]). ∎
@@
Let $t$ be a Dehn twist about the curve $a$ in $M$. We can then find a compact surface $\Sigma$ embedded in $M$ such that $a\subset\Sigma$, the inclusion of $\Sigma$ into $M$ induces a monomorphism $\operatorname{MCG}(\Sigma,\partial\Sigma)\to\Gamma(M)$, and the genus of $\Sigma$ is at least three. In particular, $t\in\operatorname{MCG}(\Sigma,\partial\Sigma)$. As $\varphi$ is compact and as $\operatorname{MCG}(\Sigma,\partial\Sigma)$ is finitely generated, there exists a compact surface $\Sigma^{\prime}$ of $M^{\prime}$ whose inclusion induces a monomorphism $\operatorname{MCG}(\Sigma^{\prime},\partial\Sigma^{\prime})\to\Gamma(M^{\prime})$ such that $\operatorname{MCG}(\Sigma^{\prime},\partial\Sigma^{\prime})$ contains $\varphi(\operatorname{MCG}(\Sigma,\partial\Sigma))$. We can then apply a result of Bridson [10, Remark 3.1(3)] to $\varphi|_{\operatorname{MCG}(\Sigma,\partial\Sigma)}\colon\thinspace\operatorname{MCG}(\Sigma,\partial\Sigma)\to\operatorname{MCG}(\Sigma^{\prime},\partial\Sigma^{\prime})$ to see that $\varphi(t)$ is a root of a multitwist. ∎
@@
By Lemma 4.5, there exists a locally finite non-separating infinite multicurve $\{a,b,c_{1},c_{2},\ldots\}$. Let $A=\varphi_{*}(a)$, $B=\varphi_{*}(b)$, and $C_{n}=\varphi_{*}(c_{n})$ for each $n\in\mathbb{N}$. By the change of coordinates principle, for each $n\in\mathbb{N}$, there exists $f_{n}\in\Gamma_{c}(M)$ such that $f_{n}(a)=a$ and $f_{n}(b)=c_{n}$. Let $m\in\mathbb{N}$ such that $\varphi(t_{a}^{m})$ is a multitwist. It follows that $\varphi(t^{m})$ is a multitwist for any non-separating Dehn twist $t$. As $f_{n}$ commutes with $t_{a}^{m}$ and conjugates $t_{b}^{m}$ to $t_{c_{n}}^{m}$, it follows from Lemma 2.2 that $\varphi(f_{n})(A)=A$ and $\varphi(f_{n})(B)=C_{n}$. Moreover,
$\varphi(f_{n})(A\cap B)=\varphi(f_{n})(A)\cap\varphi(f_{n})(B)=A\cap C_{n},$
@@
Let us assume that $A\cap B$ is non-empty, so that $A\cap C_{n}$ is non-empty for all $n\in\mathbb{N}$. As the power set of $A$ is finite, by passing to a subsequence, we may assume that $A\cap C_{k}=A\cap C_{j}$ for all $j,k\in\mathbb{N}$. This allows us to define $X=\varphi(f_{k})(A\cap B)$ to be independent of $k$. Now observe that given $j\neq k\in\mathbb{N}$, there exists $g_{j,k}\in\Gamma(M)$ such that $g_{j,k}(a)=c_{j}$ and $g_{j,k}(b)=c_{k}$. It follows that $\varphi(g_{j,k})(A\cap B)=C_{j}\cap C_{k}$, and hence $|A\cap B|=|C_{j}\cap C_{k}|$. As $|X|=|A\cap B|$ and as $X\subset C_{j}\cap C_{k}$, we have that $C_{j}\cap C_{k}=X$.
@@
For $k\in\mathbb{N}$, let $t_{k}$ be the Dehn twist about $c_{k}$. Let $m\in\mathbb{N}$ be as above, so that $\varphi(t_{k}^{m})$ is a multitwist, denoted $T_{k}$, about the multicurve $C_{k}$. As $X\subset C_{k}$, there exist multitwists $T_{X,k}$ and $T^{\prime}_{k}$ about $X$ and $C_{k}\smallsetminus X$, respectively, such that $T_{k}=T_{X,k}T^{\prime}_{k}$. Choose $h_{k}\in\Gamma_{c}(M)$ so that $h_{k}(c_{1})=c_{k}$. It follows that $\varphi(h_{k})$ conjugates $T_{1}$ to $T_{k}$. From the previous paragraph, we know that $\varphi(h_{k})$ must fix $X$ (setwise), and so $\varphi(h_{k})$ conjugates $T_{X,1}$ to $T_{X,k}$. Since $X$ is finite, its permutation group is finite; hence, by passing to a subsequence, we may assume that $\varphi(h_{k})$ induces the same permutation on $X$ for all $k$. Therefore, we have $T_{X,i}=T_{X,j}$ for all $i,j\in\mathbb{N}$. This allows us to define $T_{X}=T_{X,1}$ and to write
@@
From the fact that the multicurve $\{c_{k}\}_{k\in\mathbb{N}}$ is locally finite, the limit of the $F_{n}$ exists; however, the sequence $\{\varphi(F_{n})\}_{n\in\mathbb{N}}$ diverges. This is not possible as $\varphi$ is continuous, and therefore, $A\cap B$ must be empty. ∎
@@
From the above, we observe that if each component of $M_{a}^{\prime}$ has infinite type, then $\varphi(t_{a})$ is a multitwist. It then follows that $\varphi(t_{b})$ is also a multitwist, and as $A^{\prime}$ and $B^{\prime}$ are disjoint multicurves, $\varphi(t_{a})$ and $\varphi(t_{b})$ have disjoint supports, as desired.
We may therefore assume that $M_{a}^{\prime}$ contains at least one finite-type component. Let $F_{a}^{\prime}$ be the closures in $M^{\prime}$ of the finite-type components of $M_{a}^{\prime}$. Then there exists a representative of $\varphi(t_{a})$ supported in the union of $F_{a}^{\prime}$ with annuli about the curves in $A^{\prime}$.
@@
Every surface of genus at least three contains an embedded 4-holed sphere $\Sigma$ in which each of its boundary components and each curve contained in $\Sigma$ is non-separating in the ambient surface. Let $a_{1}$, $a_{2}$, $a_{3}$, and $a_{4}$ be the boundary components of $\Sigma$. By the lantern relation (see [14, Proposition 5.1]), there exist curves $a_{5}$, $a_{6}$, and $a_{7}$ in $\Sigma$ such that

### v2

@@ 1 Introduction
A fundamental tool in studying mapping class groups of finite-type surfaces is the forgetful homomorphism: Given a non-compact finite-type surface $S$ and a point $p\in S$, there is a surjective homomorphism $\operatorname{PMCG}(S\smallsetminus\{p\})\to\operatorname{PMCG}(S)$ obtained by “forgetting the puncture pp”, where $\operatorname{PMCG}(S)$ is the pure mapping class group. For the forgetful homomorphism, the surface in the codomain has one less end than the surface in the domain. The motivating question of this article is to ask if the same phenomenon can occur in the setting of infinite-genus surfaces without planar ends, or in other words,
@@
Let $M$ and $M^{\prime}$ be orientable 2-manifolds, and suppose $M$ has infinite genus and has no planar ends. If $G$ is a mostly pure large subgroup of $\operatorname{MCG}(M)$ and $G^{\prime}$ is a large subgroup of $\operatorname{MCG}(M^{\prime})$, then every continuous epimorphism $G\to G^{\prime}$ is induced by a homeomorphism.
@@
Next, recall that an object $A$ in a category is Hopfian if every epimorphism $A\to A$ is an automorphism. It is a standard fact that every finitely generated residually finite group is Hopfian (in the category of abstract groups). As the mapping class group of a finite-type surface is residually finite (see [15, §6.4]) and finitely generated (see [15, §4.3]), every finitely generated subgroup of the mapping class group of a finite-type orientable 2-manifold is Hopfian. One of the motivating problems for this article is to understand when $\operatorname{MCG}(M)$ and $\operatorname{PMCG}(M)$ are Hopfian groups. As a corollary of Theorem 5.1, we establish a class of big mapping class groups that are Hopfian in the category of topological groups, see Problem 4.5 of [1].
@@
A 2-manifold $M$ is perfectly self-similar if M#MM\#M is homeomorphic to $M$ and for every proper compact subset $K$ of $M$ there exists $f\in\operatorname{Homeo}(M)$ such that $f(K)\cap K=\varnothing$.
@@
In Section 5.1, in addition to proving Theorem 5.5, we extend the theorem to a larger subclass of self-similar 2-manifolds using [10], as well as give an example of a 2-manifold for which the proof techniques do not apply.
As a corollary to Theorem 5.5, we have the Hopfian property.
@@ Acknowledgements
The authors thank Greg Conner for a helpful discussion about infinitely multiplicative homomorphisms, and we thank the anonymous referee for their careful reading and helpful suggestions, which undoubtedly improved the exposition.
@@ 2.2 Mapping class groups and their topology
It will be useful to have a more workable description of this topology; this involves using a corollary to the Alexander method (see [15, §2.3] for finite-type surfaces and see [16] or [20] for infinite-type surfaces).
@@ 2.3 Twists
We already established the backwards direction above, so we only need to show the forwards direction. We will argue the contrapositive: let $A$ fail to be locally finite. Then there exists a curve $c$ and a subsequence of $\{a_{n_{j}}\}$ such that $i(c,a_{n_{j}})\neq 0$ for all $j\in\mathbb{N}$. Let $T_{m}=\prod_{j=1}^{n_{m}}t_{a_{j}}^{k_{j}}$, and let $F_{m}=T_{m}^{-1}T_{m+1}$. Note that if $\lim_{n\to\infty}\prod_{j=1}^{n}t_{a_{n}}^{k_{n}}$ exists, then the sequence $\{T_{m}\}_{m\in\mathbb{N}}$ converges, and hence $\{F_{m}\}_{m\in\mathbb{N}}$ converges to the identity. We will show that the sequence $\{F_{m}\}$ does not converge to the identity, which will finish the proof.
Observe that $F_{m}(c)=t_{a_{n_{m+1}}}^{k_{n_{m+1}}}(c)$. Then, by [15, Proposition 3.2],
$i(F_{m}(c),c)=i\left(t_{a_{n_{m+1}}}^{k_{n_{m+1}}}(c),c\right)=|k_{n_{m+1}}|i(a_{n_{m+1}},c)^{2}$
and hence $i(F_{m}(c),c)>0$. It follows that $F_{m}(c)\neq c$ for all $m\in\mathbb{N}$, and in particular, there exists a neighborhood of the identity in $\operatorname{MCG}(S)$ that the sequence $\{F_{m}\}_{m\in\mathbb{N}}$ never enters, namely $\{f\in\operatorname{MCG}(S):f(c)=c\}$. Therefore, the sequence $\{F_{m}\}$ does not converge to the identity. ∎
@@ 3 Generating compactly supported mapping classes
Let us begin by recalling the Humphries generating set (see [15, §4.4.3 and §4.4.4]). Humphries’s theorem tells us that the Dehn twists about the curves shown in Figure 1 generate $\operatorname{MCG}(S,\partial S)$; in particular, for a compact surface with nonempty boundary, the Humphries generating set consists of $2g+b$ non-separating Dehn twists (the figure shows a genus three surface with three boundary components, but it should be clear how to generalize to higher genus and more boundary components). We note that Humphries’s theorem requires the surface to be orientable and nonplanar. We will build on Humphries’s theorem to construct a generating set for $\Gamma_{c}(M)$ when $M$ is an orientable infinite-genus 2-manifold without planar ends.
@@
In what follows, we will require a generating set for the mapping class group of a four-holed sphere. Let $R$ be a compact four-holed sphere. If $a$ and $b$ are curves with $i(a,b)=2$, then $t_{a}$ and $t_{b}$ together with the Dehn twists along any three of the four boundary components of $R$ generate $\operatorname{MCG}(R,\partial R)$ (this is an exercise combining the computation of generators for the mapping class group of a pair of pants [15, §3.6.4] and a four-punctured sphere [15, §4.2.4], the inclusion homomorphism [15, Proposition 3.18], and the lantern relation [15, Proposition 5.1]).
Given a compact subsurface $\Sigma$ of a 2-manifold $M$ such that each component of $\partial\Sigma$ is separating and essential, the inclusion of $\Sigma$ into $M$ induces a monomorphism $\operatorname{MCG}(\Sigma,\partial\Sigma)\to\operatorname{MCG}(M)$, see [15, Theorem 3.18]. Abusing notation, we will identify $\operatorname{MCG}(\Sigma,\partial\Sigma)$ with its image.
Figure 2: The Alexander chain constructed in Theorem 3.2 on a 2-ended infinite-genus 2-manifold.
@@
Let $\{\Sigma_{n}\}_{n\in\mathbb{N}}$ be the compact exhaustion of $M$ given by Lemma 3.1. Observe that
$\Gamma_{c}(M)=\bigcup_{n\in\mathbb{N}}\operatorname{MCG}(\Sigma_{n},\partial\Sigma_{n}).$
We therefore need to give an Alexander chain $\mathcal{A}$ such that the Dehn twists about the curves in $\mathcal{A}$ contained in $\Sigma_{n}$ generate $\operatorname{MCG}(\Sigma_{n},\partial\Sigma_{n})$. Let us construct $\mathcal{A}$ and then argue that it satisfies this property.
@@
Let $S_{0}=\Sigma_{0}$, and for $n\in\mathbb{N}$, let $S_{n}$ be the closure of $\Sigma_{n}\smallsetminus\Sigma_{n-1}$. Let $\mathcal{A}_{0}$ consist of any two simple closed curves with geometric intersection number one; hence, the curves in $\mathcal{A}_{0}$ generate $\operatorname{MCG}(\Sigma_{0},\partial\Sigma_{0})$ (see [15, §3.6.4]). For $i\in\{2,3\}$, let $F_{i}$ be the genus one compact surface with $i$ boundary components. Let $C_{i}$ be the Alexander chain in $F_{i}$ shown in Figure 3. For $n\in\mathbb{N}$, let $i(n)\in\{2,3\}$ such that $S_{n}$ is homeomorphic to $F_{i(n)}$, and fix a homeomorphism $f_{n}\colon\thinspace F_{i(n)}\to S_{n}$. Let $\mathcal{A}^{\prime}$ be the chain defined by $\mathcal{A}^{\prime}=\mathcal{A}_{0}\cup\bigcup_{n=1}^{\infty}f_{n}(C_{i(n)})$.
We now add curves to the collection $\mathcal{A}^{\prime}$ so that it forms an Alexander chain. Each of the curves in Figure 3 are colored either blue or red; each surface in the figure contains a single red curve that intersects each of the remaining curves—the blue curves—exactly once. Note that, in each surface, the blue curves give a pants decomposition. Designate one curve in $\mathcal{A}_{0}$ to be blue. Now, let $\mathcal{A}^{\prime}_{b}$ denote the subset of $\mathcal{A}^{\prime}$ containing all the blue curves. Each component of $M\smallsetminus\mathcal{A}^{\prime}_{b}$ is a four-holed sphere. Enumerate the components as follows: define $\Omega_{n}$ to be the unique component that is contained in $\Sigma_{n}$ but not in $\Sigma_{n-1}$. Choose a non-separating curve $a_{n}$ in $M$ that is contained in $\Omega_{n}$ and such that $\mathcal{A}^{\prime}\cup\{a_{n}\}$ is a chain. Let $\mathcal{A}=\mathcal{A}^{\prime}\cup\{a_{n}:n\in\mathbb{N}\}$. By construction, $\mathcal{A}$ is an Alexander chain. An example of the chain $\mathcal{A}$ can be seen in Figure 2.
@@ 4 Irreducible epimorphisms
Using our assumptions and [11], it is not hard to see that $\varphi$ must map a Dehn twist $t_{a}$ about a curve $a$ to a root of a Dehn twist about a multicurve $\varphi_{*}(a)$ (Lemma 4.4). We then eliminate roots via a sequence of steps, taking advantage of the infinite-type setting. The first step (Lemma 4.6) shows that if $a$ and $b$ are disjoint and non-separating (individually and jointly) then $\varphi_{*}(a)$ and $\varphi_{*}(b)$ are also disjoint in a strong sense (see definition below). This result relies on a pigeon-hole argument tailored to the infinite-type setting, based on the existence of a locally finite, non-separating, infinite multicurve containing the curves $a$ and $b$ (Lemma 4.5). Next, a continuity argument using an infinite-product trick shows that local finiteness of multicurves is preserved under $\varphi_{*}$ (Lemma 4.7). This gives rise to a second pigeon-hole argument that allows us to promote disjointness to the supports of the roots themselves (Lemma 4.8). Once the disjointness of supports is established, we can use the lantern relation to eliminate roots (Lemma 4.9). Finally, we eliminate multiple components in $\varphi_{*}(a)$ for a non-separating $a$ (Proposition 4.12), which also shows that intersection number one is preserved under $\varphi$. This part of the argument relies on the recent work of [13] on the braid relation between multitwists. From here, we move to the proof of Theorem 4.1, where we use the results of [6] to obtain homeomorphisms between compact exhaustions of $M$ and $M^{\prime}$, allowing us to take a direct limit to construct the desired homeomorphism $M\to M^{\prime}$.
@@
If $g\in\Gamma_{c}(M)$ is nontrivial, then there exists a finite-type subsurface $\Sigma\subset M$ containing the support of $g$ such that each boundary component of $\Sigma$ is an essential simple closed curve. In this case, the inclusion $\Sigma\hookrightarrow M$ induces a monomorphism $\operatorname{MCG}(\Sigma,\partial\Sigma)\to\Gamma_{c}(M)$ whose image contains $g$. Therefore, we can view $g$ as an element of $\operatorname{MCG}(\Sigma,\partial\Sigma)$, which is torsion free (see [15, Corollary 7.3]). ∎
@@
Let $t$ be a Dehn twist about the curve $a$ in $M$. We can then find a compact surface $\Sigma$ embedded in $M$ such that $a\subset\Sigma$, the inclusion of $\Sigma$ into $M$ induces a monomorphism $\operatorname{MCG}(\Sigma,\partial\Sigma)\to\Gamma(M)$, and the genus of $\Sigma$ is at least three. In particular, $t\in\operatorname{MCG}(\Sigma,\partial\Sigma)$. As $\varphi$ is compact and as $\operatorname{MCG}(\Sigma,\partial\Sigma)$ is finitely generated, there exists a compact surface $\Sigma^{\prime}$ of $M^{\prime}$ whose inclusion induces a monomorphism $\operatorname{MCG}(\Sigma^{\prime},\partial\Sigma^{\prime})\to\Gamma(M^{\prime})$ such that $\operatorname{MCG}(\Sigma^{\prime},\partial\Sigma^{\prime})$ contains $\varphi(\operatorname{MCG}(\Sigma,\partial\Sigma))$. We can then apply a result of Bridson [11, Remark 3.1(3)] to $\varphi|_{\operatorname{MCG}(\Sigma,\partial\Sigma)}\colon\thinspace\operatorname{MCG}(\Sigma,\partial\Sigma)\to\operatorname{MCG}(\Sigma^{\prime},\partial\Sigma^{\prime})$ to see that $\varphi(t)$ is a root of a multitwist. ∎
@@
By Lemma 4.5, there exists a locally finite non-separating infinite multicurve $\{a,b,c_{1},c_{2},\ldots\}$. Let $A^{\prime}=\varphi_{*}(a)$, $B^{\prime}=\varphi_{*}(b)$, and $C^{\prime}_{n}=\varphi_{*}(c_{n})$ for each $n\in\mathbb{N}$. By the change of coordinates principle, for each $n\in\mathbb{N}$, there exists $f_{n}\in\Gamma_{c}(M)$ such that $f_{n}(a)=a$ and $f_{n}(b)=c_{n}$. Let $m\in\mathbb{N}$ such that $\varphi(t_{a}^{m})$ is a multitwist. It follows that $\varphi(t^{m})$ is a multitwist for any non-separating Dehn twist $t$. As $f_{n}$ commutes with $t_{a}^{m}$ and conjugates $t_{b}^{m}$ to $t_{c_{n}}^{m}$, it follows from Lemma 2.2 that $\varphi(f_{n})(A^{\prime})=A^{\prime}$ and $\varphi(f_{n})(B^{\prime})=C^{\prime}_{n}$. Moreover,
$\varphi(f_{n})(A^{\prime}\cap B^{\prime})=\varphi(f_{n})(A^{\prime})\cap\varphi(f_{n})(B^{\prime})=A^{\prime}\cap C^{\prime}_{n},$
@@
Let us assume that $A^{\prime}\cap B^{\prime}$ is non-empty, so that $A^{\prime}\cap C^{\prime}_{n}$ is non-empty for all $n\in\mathbb{N}$. As the power set of $A^{\prime}$ is finite, by passing to a subsequence, we may assume that $A^{\prime}\cap C^{\prime}_{k}=A^{\prime}\cap C^{\prime}_{j}$ for all $j,k\in\mathbb{N}$. This allows us to define $X=\varphi(f_{k})(A^{\prime}\cap B^{\prime})$ to be independent of $k$. Now observe that given $j\neq k\in\mathbb{N}$, there exists $g_{j,k}\in\Gamma(M)$ such that $g_{j,k}(a)=c_{j}$ and $g_{j,k}(b)=c_{k}$. It follows that $\varphi(g_{j,k})(A^{\prime}\cap B^{\prime})=C^{\prime}_{j}\cap C^{\prime}_{k}$, and hence $|A^{\prime}\cap B^{\prime}|=|C^{\prime}_{j}\cap C^{\prime}_{k}|$. As $|X|=|A^{\prime}\cap B^{\prime}|$ and as $X\subset C^{\prime}_{j}\cap C^{\prime}_{k}$, we have that $C^{\prime}_{j}\cap C^{\prime}_{k}=X$.
@@
For $k\in\mathbb{N}$, let $t_{k}$ be the Dehn twist about $c_{k}$. Let $m\in\mathbb{N}$ be as above, so that $\varphi(t_{k}^{m})$ is a multitwist, denoted $T_{k}$, about the multicurve $C^{\prime}_{k}$. As $X\subset C^{\prime}_{k}$, there exist multitwists $T_{X,k}$ and $T^{\prime}_{k}$ about $X$ and $C^{\prime}_{k}\smallsetminus X$, respectively, such that $T_{k}=T_{X,k}T^{\prime}_{k}$. Choose $h_{k}\in\Gamma_{c}(M)$ so that $h_{k}(c_{1})=c_{k}$. It follows that $\varphi(h_{k})$ conjugates $T_{1}$ to $T_{k}$. From the previous paragraph, we know that $\varphi(h_{k})$ must fix $X$ (setwise), and so $\varphi(h_{k})$ conjugates $T_{X,1}$ to $T_{X,k}$. Since $X$ is finite, its permutation group is finite; hence, by passing to a subsequence, we may assume that $\varphi(h_{k})$ induces the same permutation on $X$ for all $k$. Therefore, we have $T_{X,i}=T_{X,j}$ for all $i,j\in\mathbb{N}$. This allows us to define $T_{X}=T_{X,1}$ and to write
@@
From the fact that the multicurve $\{c_{k}\}_{k\in\mathbb{N}}$ is locally finite, the limit of the $F_{n}$ exists; however, the sequence $\{\varphi(F_{n})\}_{n\in\mathbb{N}}$ diverges. This is not possible as $\varphi$ is continuous, and therefore, $A^{\prime}\cap B^{\prime}$ must be empty. ∎
@@
From the above, we observe that if each component of $M_{a}^{\prime}$ has infinite type, then $\varphi(t_{a})$ is a multitwist. As $\varphi(t_{b})$ is conjugate to $\varphi(t_{a})$, it is also a multitwist, and as $A^{\prime}$ and $B^{\prime}$ are disjoint multicurves, $\varphi(t_{a})$ and $\varphi(t_{b})$ have disjoint supports, as desired.
The previous paragraph dealt with the case in which $M_{a}^{\prime}$ has no finite-type component, so we may now assume that $M_{a}^{\prime}$ contains at least one finite-type component. Let $F_{a}^{\prime}$ be the closures in $M^{\prime}$ of the finite-type components of $M_{a}^{\prime}$. Then there exists a representative of $\varphi(t_{a})$ supported in the union of $F_{a}^{\prime}$ with annuli about the curves in $A^{\prime}$.
@@
Every surface of genus at least three contains an embedded 4-holed sphere $\Sigma$ in which each of its boundary components and each curve contained in $\Sigma$ is non-separating in the ambient surface. Let $a_{1}$, $a_{2}$, $a_{3}$, and $a_{4}$ be the boundary components of $\Sigma$. By the lantern relation (see [15, Proposition 5.1]), there exist curves $a_{5}$, $a_{6}$, and $a_{7}$ in $\Sigma$ such that